a boatman can row his boat at a speed of 8km per hour in still water.if a river flows at a speed of 2km per hour,then how long will the boatman take to row his boat 1200metres against the direction of the current.
Answer by TimothyLamb(4379)

s = 8 - 2 = 6 kph
---
s = d/t
t = d/s
---
t = 1.2 km / 6 kph = 0.2 hours * 60 min/hour = 12 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php